The velocity of money in the small Republic of Sloagia is always the same. Last year, the money supply was $ 5 billion and real GDP was $ 20 billion.

This year, the money supply increased by 5 percent, real GDP by 4.5 percent, and nominal GDP is $ 19 billion.

    

           Calculate the velocity of money  

           The price level last year=

           The Price level this year =

           The inflation rate=
Answer #1

Answer:

Velocity of money = 3.62

Price level last year = 0.905

Price level this year = 0.909

Inflation rate = 0.44%

Calculation:

MV = PY

(PY = Nominal GDP)

MV = Nominal GDP

PY Velocity of money(V) =M Nominal GDP Money supply

Last year,

P* 20 V =-

V = 4P

----------------------------------------------------------------------------------------------------------------------

This year, money supply increases by 5%.

Money supply = $5 + ($5 * 5%) = $5 + $0.25

Money supply = $5.25

Nominal GDP = $19 (Given)

Nominal GDP Velocity of money(V) = ? Money supply

Velocity of money(V)= 15 = 3.62 _1920

-----------------------------------------------------------------------------------------------------------------------

Given assumption: Velocity of money is always the same.

Price level last year :

V = 4P

3.62 = 4P

P = 3.62/4

P = 0.905

-----------------------------------------------------------------------------------------------------------------------

Price level in this year:

MV = PY

MV P = V

Real GDP (Y) increases by 4.5% = $20 + ($20*4.5%) = $20 + 0.9 = $20.9

P = 5.25 * 3.62 -= 0.909 20,9

-----------------------------------------------------------------------------------------------------------------------

Inflation rate = Price level this year - Price level last year Price level last year -* 100

0.004 Inflation rate = 0.909 - 0.905 -* 100 = 0.905 0.905 * 100

Inflation rate = 0.44%